Find the future value, using the future value formula and a calculator. (Round your answer to the nearest cent.)
$889 at 5.5% compounded quarterly for 5 years

Answer by rfer(16322)

FV=889(1+0.055/4)^(4*5)
FV=889(1.01375)^20
FV=889(1.3141)
FV=1168.21